OpenStudy (anonymous):

A differentiable function f has the property that f(5)=3 and fprime(5)=4. What is the estimate for f(4.8) using the local linear approximation for f at x=5?

OpenStudy (anonymous):

The linear approximation means that you assume the function is linear. So just do f(5) +f'(5) (x-5)

OpenStudy (anonymous):

This works out, in this specific case to give me: 3 + 4(-0.2)

OpenStudy (anonymous):

thanks i understand everything except for the 0.2 part

OpenStudy (anonymous):

how do you know to do 4.8 - 5?

OpenStudy (anonymous):

For a linear function, I have something like y = mx + b. m is the CONSTANT slope. If I increase x by some amount, and I want to know how much y has increased, I just multiply m by the increase in x. So, let's say the slope is 2, and x goes from 4 to 7, then x has increased by 3, which means y will have increased by m*x = 2*3 = 6 Now, if x goes from 8 to 7.5, then x has increased by (-0.5), so y will have increased by m * x = 2(-0.5) = -1, which means y went down by 1.

OpenStudy (anonymous):

thank you you so much!

OpenStudy (anonymous):

You're welcome =D